OpenStudy (anonymous):

Prove , cos20cos40cos80=1/8 . Angles are in degrees .

OpenStudy (anonymous):

blue?

OpenStudy (anonymous):

Blue what?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

blue is the answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What the??

OpenStudy (anonymous):

Let cos20*cos40*cos80 = U. Therefore 8*U*sin20 = 4*(2*sin20*cos20)*cos40*cos80 = 4*sin(2*20)*cos40*cos80 [since sin2A = 2*sinA*cosA] = 2*(2*sin40*cos40)*cos80 = 2*sin(2*40)*cos80 = 2*sin80*cos80 = sin(2*80) = sin160 = sin(180-20) = sin20 [since sin(180-x) = sin(x)] So U = (sin20)/(8*sin20) = 1/8

OpenStudy (anonymous):

calm down Erkot:)

OpenStudy (anonymous):

Grammatical mistake . My bad . Thank you !

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

guess where i am blue?

OpenStudy (anonymous):

lab?